OpenStudy (anonymous):

major math problem...help!!! find the vertex, focus, and directrix of the paarabola and sketch it graph. (x+3/2)^2=4(y-2) i have found the vertex and it is (-3/2,2)... i'm not sure how to find the others :(

sam (.sam.):

rearrange the equation (x+3/2)^2=4(y-2) y=(x+3/2)^2/4+2

OpenStudy (anonymous):

can u show step-by-step how u did that plz?

sam (.sam.):

rearrange the equation (x+3/2)^2=4(y-2) divide 4 both sides rearrange the equation (x+3/2)^2/4=(y-2) add 2 to both sides y=(x+3/2)^2/4+2

sam (.sam.):

then expand (x+3/2)^2, \[y=\frac{(x+3/2)^2}{4}+2\] \[y=\frac{x^2+3x+\frac{9}{4}}{4}+2\] multiply 4 \[y=\frac{4x^2+12x+9}{16}+2\] \[y=\frac{x^2}{4}+\frac{3 x}{4}+\frac{41}{16}\]

OpenStudy (anonymous):

what is that above the 4 in the second equation???

sam (.sam.):

9

OpenStudy (anonymous):

o. :)

OpenStudy (anonymous):

whats the next step?

OpenStudy (anonymous):

the vertex i had was correct right?

sam (.sam.):

yes

OpenStudy (anonymous):

okay... then how could get the focus for this parabola?

sam (.sam.):

use \[y=\frac{x^2}{4}+\frac{3 x}{4}+\frac{41}{16}\] find the y-intercept

OpenStudy (anonymous):

so i would set x to zero right?

sam (.sam.):

yes

OpenStudy (anonymous):

okay let me do it and u could tell me if im right

OpenStudy (anonymous):

41/16?

sam (.sam.):

yes

OpenStudy (anonymous):

okay so the my focus is (-3/2, 41/16?

sam (.sam.):

focus is \[\left(-\frac{3}{2},3\right)\]

OpenStudy (anonymous):

how is it 3???????

sam (.sam.):

the complete the square form of \[y=\frac{x^2}{4}+\frac{3 x}{4}+\frac{41}{16}\] \[\frac{1}{4} \left(x+\frac{3}{2}\right)^2+2\] p=1/4(1/4)=1 focus=( h, c+1) focus=(-3/2 , 2+p) focus=(-3/2 ,3)

OpenStudy (anonymous):

?

sam (.sam.):

?

OpenStudy (anonymous):

lol no i need help w/ what u just said or wrote

sam (.sam.):

\[\frac{1}{4} \left(x+\frac{3}{2}\right)^2+2\] \[a \left(x+b\right)^2+c\] \[p=\frac{1}{4a}\] \[focus=( b, c+p) \]

OpenStudy (anonymous):

oka so that equation is used for what?

sam (.sam.):

you mean p?

OpenStudy (anonymous):

yeas

sam (.sam.):

that's a formula you need to memorize

OpenStudy (anonymous):

and what exactly does that formula solve for?

sam (.sam.):

focus

OpenStudy (anonymous):

ooooooooooooo really??? so that the formula to find the focus on all problem right?